9- How much time is required to administer 0.500 L of glucose 5% IV solution if the flow rate is 35 mL/h?
10-A baby weighs 8 lb and 4 oz. Calculate the weight in grams
Answer #1

Time taken = Volume needed / rate of delivery = 500 / 35 mL/h = 14.286 hr

1 oz = 0.0625 lb

4 oz = 4 * 0.0625 lb = 0.375 lb

Mass of baby = 8.375 lb

Mass in kg = mass in lb * 0.454 = 8.375 * 0.454 = 3.802 kg = 380.2 g